# Approximation by perfect complexes detects Rouquier dimension

### v1

@@ 3. Approximations
Assume Setup 3.21. For an object $K$ of $D^{b}_{\operatorname{coh}}(X)$, there is a canonical isomorphism
@@
where the colimit is over Diagram 3.0.1. This isomorphism is defined as follows. The $n^{th}$ map
@@
(3.0.2) $G\xrightarrow{1/b^{n}}G\to\mathbb{L}f^{\ast}\mathbb{R}g_{\ast}\overline{G}\xrightarrow{\mathbb{L}f^{\ast}(\varphi)}\mathbb{L}f^{\ast}K,$
@@
Assume Setup 3.21. If $\psi\colon K\to L$ is an $\mathbb{R}g_{\ast}\overline{G}$-ghost map in $D^{b}_{\operatorname{coh}}(X)$, then $\mathbb{L}f^{\ast}(\psi)$ is a $G$-ghost map in $D^{b}_{\operatorname{coh}}(U)$.
@@
Assume $\mathbb{L}f^{\ast}(\psi)\colon\mathbb{L}f^{\ast}K\to\mathbb{L}f^{\ast}L$ is not a $G$-ghost map. After replacing, if necessary, $\psi$ by some shift $\psi[i]$, this means there is a nonzero composition $G\to\mathbb{L}f^{\ast}K\to\mathbb{L}f^{\ast}L$. By Lemma 3.22, there exist an integer $n$ and a map $\varphi\colon\mathbb{R}g_{\ast}\overline{G}\to K$ such that $G\to\mathbb{L}f^{\ast}K$ is equal to a composition of the form in Diagram 3.0.2. But then from this diagram and the fact that $G\to\mathbb{L}f^{\ast}K\to\mathbb{L}f^{\ast}L$ is nonzero, we see that $\mathbb{L}f^{\ast}(\psi)\circ\mathbb{L}f^{\ast}(\varphi)\neq 0$, and hence $\psi\circ\varphi\neq 0$. This tells us $\psi$ is not a $\mathbb{R}g_{\ast}\overline{G}$-ghost map. ∎
@@
Let $f\colon U\to X$ be an étale morphism of Noetherian affine schemes. Given an object $G$ of $D^{b}_{\operatorname{coh}}(U)$, there exists an object $H$ of $D^{b}_{\operatorname{coh}}(X)$ with the following property: If $\psi\colon K\to L$ is an $H$-ghost map in $D^{b}_{\operatorname{coh}}(X)$, then $\mathbb{L}f^{\ast}(\psi)$ is a $G$-ghost map in $D^{b}_{\operatorname{coh}}(X)$.
@@
[…]
@@
We will show this more generally for any local homomorphism $R\to S$ of Noetherian local rings which is a filtered colimit $S=\operatorname{colim}_{i}S_{i}$ of étale ring maps $R\to S_{i}$. To show the inequality, we will show every $K\in D^{b}_{\operatorname{coh}}(S)$ is a direct summand of $L\otimes_{R}^{\mathbb{L}}S$ for some object $L\in D^{b}_{\operatorname{coh}}(R)$. So let $K\in D^{b}_{\operatorname{coh}}(S)$. Then by Lemma 3.26, there exists an $i$ and on object $K_{i}\in D^{b}_{\operatorname{coh}}(S_{i})$ such that $K_{i}\otimes_{S_{i}}^{\mathbb{L}}S\cong K$, and then by Theorem 1.2 there is an object $L\in D^{b}_{\operatorname{coh}}(R)$ such that $K_{i}$ is a direct summand of $L\otimes_{R}^{\mathbb{L}}S_{i}$. But then $K$ is a direct summand of $L\otimes_{R}^{\mathbb{L}}S$.
Now assume $R$ admits a dualizing complex. We will show the reverse inequality. Let $d=\operatorname{dim}(D^{b}_{\operatorname{coh}}(S))$ and choose $G\in D^{b}_{\operatorname{coh}}(S)$ such that $D^{b}_{\operatorname{coh}}(S)=\langle G\rangle_{d+1}$. Then by Lemma 3.26 again, there exists an $i$ and an object $G_{i}\in D^{b}_{\operatorname{coh}}(S_{i})$ such that $G_{i}\otimes_{S_{i}}^{\mathbb{L}}S\cong G$. By Lemma 3.25, there exists an object $H\in D^{b}_{\operatorname{coh}}(R)$ such the base change of an $H$-ghost along $R\to S_{i}$ is a $G_{i}$-ghost. But then by Lemma 3.24, the base change of an $H$-ghost along $R\to S$ is a $G$-ghost, and we conclude as in the proof of Theorem 1.3. ∎

### v2

@@ 3. Approximations
Assume 3.20. For an object $K$ of $D^{b}_{\operatorname{coh}}(X)$, there is a canonical isomorphism
@@
where the colimit is over Diagram 3.1. This isomorphism is defined as follows. The $n^{th}$ map
@@
(3.2) $G\xrightarrow{1/b^{n}}G\to\mathbb{L}f^{\ast}\mathbb{R}g_{\ast}\overline{G}\xrightarrow{\mathbb{L}f^{\ast}(\varphi)}\mathbb{L}f^{\ast}K,$
@@
Assume 3.20. If $\psi\colon K\to L$ is an $\mathbb{R}g_{\ast}\overline{G}$-ghost map in $D^{b}_{\operatorname{coh}}(X)$, then $\mathbb{L}f^{\ast}(\psi)$ is a $G$-ghost map in $D^{b}_{\operatorname{coh}}(U)$.
@@
Assume $\mathbb{L}f^{\ast}(\psi)\colon\mathbb{L}f^{\ast}K\to\mathbb{L}f^{\ast}L$ is not a $G$-ghost map. After replacing, if necessary, $\psi$ by some shift $\psi[i]$, this means there is a nonzero composition $G\to\mathbb{L}f^{\ast}K\to\mathbb{L}f^{\ast}L$. By Lemma 3.21, there exist an integer $n$ and a map $\varphi\colon\mathbb{R}g_{\ast}\overline{G}\to K$ such that $G\to\mathbb{L}f^{\ast}K$ is equal to a composition of the form in Diagram 3.2. But then from this diagram and the fact that $G\to\mathbb{L}f^{\ast}K\to\mathbb{L}f^{\ast}L$ is nonzero, we see that $\mathbb{L}f^{\ast}(\psi)\circ\mathbb{L}f^{\ast}(\varphi)\neq 0$, and hence $\psi\circ\varphi\neq 0$. This tells us $\psi$ is not a $\mathbb{R}g_{\ast}\overline{G}$-ghost map. ∎
@@
Let $f\colon U\to X$ be an étale morphism of Noetherian affine schemes. Given an object $G$ of $D^{b}_{\operatorname{coh}}(U)$, there exists an object $H$ of $D^{b}_{\operatorname{coh}}(X)$ with the following property: If $\psi\colon K\to L$ is an $H$-ghost map in $D^{b}_{\operatorname{coh}}(X)$, then $\mathbb{L}f^{\ast}(\psi)$ is a $G$-ghost map in $D^{b}_{\operatorname{coh}}(U)$.
@@
It follows from Theorem 1.2(1) that $\dim D^{b}_{\operatorname{coh}}(X)\geq\dim D^{b}_{\operatorname{coh}}(U)$, so we check the reverse inequality. Let $d=\operatorname{dim}D^{b}_{\operatorname{coh}}(U)$ and choose an object $G$ such that $\langle G\rangle_{d+1}=D^{b}_{\operatorname{coh}}(U)$. Let $H\in D^{b}_{\operatorname{coh}}(X)$ be as in Lemma 3.24. We claim $\langle H\rangle_{d+1}=D^{b}_{\operatorname{coh}}(X)$. Our assumption on dualizing complexes lets us appeal to Letz’s converse ghost lemma [Let21, §$\S 2.13$], and so, from Remark 2.10, it suffices to show a composition of $d+1$ $H$-ghosts in $D^{b}_{\operatorname{coh}}(X)$ vanishes. By Lemma 3.24 and the ghost lemma [Let21, §$\S 2.13$], the pullback of such a composition to $U$ vanishes. But then by part (2) of Lemma 3.23, the composition itself vanishes, completing the proof. ∎
@@
Here “colimit” means 2-colimit in the 2-category of categories. More explicitly, the Lemma says that every object of $D^{b}_{\operatorname{coh}}(X)$ is isomorphic to the derived pullback of an object of $D^{b}_{\operatorname{coh}}(X_{i})$ for some $i$, and if $K_{i},L_{i}\in D^{b}_{\operatorname{coh}}(X)$, then
$\operatorname{Hom}_{D^{b}_{\operatorname{coh}}(X)}(K,L)=\operatorname{colim}_{j\geq i}\operatorname{Hom}_{D^{b}_{\operatorname{coh}}(X_{j})}(K_{j},L_{j}),$
where $K,L$ (resp. $K_{j},L_{j}$) are the pullbacks of $K_{i},L_{i}$ to $X$ (resp. $X_{j}$).
@@
Note that Henselization of our Noetherian local ring remains Noetherian, see [Sta23, Tag 06LJ]. We will show this more generally for any local homomorphism $R\to S$ of Noetherian local rings which is a filtered colimit $S=\operatorname{colim}_{i}S_{i}$ of étale ring maps $R\to S_{i}$. To show the inequality, we will show every $K\in D^{b}_{\operatorname{coh}}(S)$ is a direct summand of $L\otimes_{R}^{\mathbb{L}}S$ for some object $L\in D^{b}_{\operatorname{coh}}(R)$. So let $K\in D^{b}_{\operatorname{coh}}(S)$. Then by Lemma 3.25, there exists an $i$ and on object $K_{i}\in D^{b}_{\operatorname{coh}}(S_{i})$ such that $K_{i}\otimes_{S_{i}}^{\mathbb{L}}S\cong K$, and then by Theorem 1.2 there is an object $L\in D^{b}_{\operatorname{coh}}(R)$ such that $K_{i}$ is a direct summand of $L\otimes_{R}^{\mathbb{L}}S_{i}$. But then $K$ is a direct summand of $L\otimes_{R}^{\mathbb{L}}S$.
Now assume $R$ admits a dualizing complex. This case allows us to appeal to Letz’s converse ghost lemma [Let21, §$\S 2.13$] and Remark 2.10. We will show the reverse inequality. Let $d=\operatorname{dim}(D^{b}_{\operatorname{coh}}(S))$ and choose $G\in D^{b}_{\operatorname{coh}}(S)$ such that $D^{b}_{\operatorname{coh}}(S)=\langle G\rangle_{d+1}$. Then by Lemma 3.25 again, there exists an $i$ and an object $G_{i}\in D^{b}_{\operatorname{coh}}(S_{i})$ such that $G_{i}\otimes_{S_{i}}^{\mathbb{L}}S\cong G$. By Lemma 3.24, there exists an object $H\in D^{b}_{\operatorname{coh}}(R)$ such the base change of an $H$-ghost along $R\to S_{i}$ is a $G_{i}$-ghost. But then by Lemma 3.23, the base change of an $H$-ghost along $R\to S$ is a $G$-ghost, and we conclude as in the proof of Theorem 1.3. ∎